Accelerator Update for November 12, 2008 - November 14, 2008
Wednesday Nov 12
The day shift began with Operations monitoring the Tevatron (TeV) store 6554 (206.79E30) and the Recycler stash (68.8E10), with the Antiproton Source stacking (30.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the NuMI, MiniBooNE, and Meson MTest experiments taking beam.
The Neutron Therapy Facility reported at 10:54 AM that they had completed treatment for their last patient of the day.
At 10:54 AM, Operations reported that a Linac RF station (LRF3) had tripped off twice. It reset.
At 11:30 AM, Linac experts reported that they were turning off LRF3 for a few minutes to investigate possible problems.
Linac experts reported at 1:03 PM that they were taking down LRF3 for 30 minutes to make repairs.
At 2:12 PM, Operations reported a large amount of database errors. Experts investigated and later reported that the power outage had caused the errors. Operations and Controls experts continued to investigate and fix problems.
A Pbar expert began a ten-minute study period at 3:07 PM.
Operations reported at 4 PM that the Recycler had stashed antiprotons six times during the day shift.
Operations disabled the MTest beam at 6:02 PM.
Operations reported at 11:56 PM that the Recycler had stashed eight times during the evening shift.
Thursday Nov 13
The midnight shift began with Operations monitoring the TeV store 6554 (29.58E30) and the stash (375.9E10), with the Pbar stacking (25.1E10), and with NuMI and MiniBooNE taking beam.
Operations terminated store 6554 at 12:24 AM.
Operations put the TeV into shot setup at 12:36 AM.
At 12:51 AM, Operations reported a MI corrector magnet power supply (I:H214) that was not following its reference, which was causing bad beam positions for the Pbar beam. This also interrupted beam to NuMI. Operators investigated and reported at 1:24 AM that they found and replaced a bad power supply regulator.
Pbar resumed stacking at 1:32 AM.
Operations established store 6555 at 3:08 AM with an initial luminosity of 309.6E30.
At 5:16 AM, a Pbar power supply (M:HV100) tripped off and would not reset. Operators investigated and called in a power supply expert.
Operations established beam to MTest at 6:34 AM.
A power supply expert repaired M:HV100. Pbar resumed stacking at 6:34 AM.
Operations reported at 8:02 AM that the Recycler had stashed four times during the midnight shift.
Operations reported at 1:18 PM that the 400 MeV area trips were due to phase problems with LRF4 and 5. Operators began phase tuning.
Operations put Pbar into standby mode at 1:25 PM.
At 2:12 PM, the 400 MeV area beam permit tripped off for the first of nine times in about 90 minutes.
Operations reported at 4 PM that the Recycler had stashed seven times during the day shift.
At 4:59 PM, the MeV 400 area beam permit tripped off for the first of two times in about 2 hours. Linac experts tuned.
Operations disabled the MTest beam at 6:04 PM.
Operations terminated store 6555 at 9:39 PM.
Operations put the TeV into shot setup at 9:50 PM.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Friday Nov 14
The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash (58.6E10), with Pbar stacking (21.2E10), and with NuMI and MiniBooNE taking beam.
Operations reported the state of the quench investigation at 5:55 AM. The two pictures below are TeV Tune Spectrograms. The left picture showed store 6555 at the beginning of its squeeze. The right picture showed the quenched store at the same time. A TeV expert reported that the red-orange-yellow horizontal spike at the top of the picture was the beam going coherent. He believes this might have been the result of the separator spark, but there was no other evidence to support that assumption.
Operations established beam to MTest at 6:26 AM.
Operations prepared the TeV sectors A-E for access at 6:56 AM. Cryo system experts will repair the stuck Kautzky valve at sector C45 and check the C3CV4 cryostat valve.
Operations prepared CDF for access at 7:36 AM.
Operations reported at 7:58 AM that the Recycler had stashed seven times during the midnight shift.
The TeV came out of access at 8:08 AM.
The Plans for Friday and the Weekend
The plans for today are to complete the TeV checkout and to go into shot setup whenever the checkout is completed. The plans for the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 11/14/2008 email Fermilab|
|Security, Privacy, Legal|